# Adapting Newton’s Method to Neural Networks through a Summary of Higher-Order Derivatives

### v3

@@ 3 Summarizing higher-order information
Let us consider the minimization of a loss function $\mathcal{L}:\mathbb{R}^{P}\rightarrow\mathbb{R}$ according to a variable $\bm{\theta}\in\mathbb{R}^{P}$.
@@
The order-$d$ derivative of $\mathcal{L}$ at a point $\bm{\theta}$, that we denote by $\frac{\mathrm{d}^{d}\mathcal{L}}{\mathrm{d}\bm{\theta}^{d}}(\bm{\theta})$, can be viewed as either a $d$-linear form (see Dieudonné (1960)) or as an order-$d$ tensor belonging to $\mathbb{R}^{P^{d}}$. For convenience, we will use the latter: the coefficients of the tensor $\mathbf{A}=\frac{\mathrm{d}^{d}\mathcal{L}}{\mathrm{d}\bm{\theta}^{d}}(\bm{\theta})\in\mathbb{R}^{P^{d}}$ are $A_{i_{1},\cdots,i_{d}}=\frac{\partial^{d}\mathcal{L}}{\partial\theta_{i_{1}}\cdots\partial\theta_{i_{d}}}(\bm{\theta})$, where $(i_{1},\cdots,i_{d})\in\{1,\cdots,P\}^{d}$ is a multi-index. For a tensor $\mathbf{A}\in\mathbb{R}^{P^{d}}$, we will use the following notation for tensor contraction:
$\forall(\mathbf{u}_{1},\cdots,\mathbf{u}_{d})\in\mathbb{R}^{P}\times\cdots\times\mathbb{R}^{P},\quad\mathbf{A}[\mathbf{u}^{1},\cdots,\mathbf{u}^{d}]:=\sum_{i_{1}=1}^{P}\cdots\sum_{i_{d}=1}^{P}A_{i_{1},\cdots,i_{d}}u^{1}_{i_{1}}\cdots u^{d}_{i_{d}}.$ (3)
The order-$d$ derivative $\frac{\mathrm{d}^{d}\mathcal{L}}{\mathrm{d}\bm{\theta}^{d}}(\bm{\theta})\in\mathbb{R}^{P^{d}}$ contains $P^{d}$ scalars. But, even when considering its symmetries, it is computationally too expensive to compute it exactly for $d\geq 2$ in most cases. For instance, it is not even possible to compute numerically the full Hessian of $\mathcal{L}$ according to the parameters of a small neural network, i.e., with $P=10^{5}$ and $d=2$, the Hessian contains $P^{d}=10^{10}$ scalars.
@@
At the opposite, one can obtain cheap higher-order information about $\mathcal{L}$ at $\bm{\theta}$ by considering a specific direction $\mathbf{u}\in\mathbb{R}^{P}$. The Taylor expansion of $\mathcal{L}(\bm{\theta}+\mathbf{u})$ gives:
$\mathcal{L}(\bm{\theta}+\mathbf{u})=\mathcal{L}(\bm{\theta})+\sum_{d=1}^{D}\frac{1}{d!}\frac{\mathrm{d}^{d}\mathcal{L}}{\mathrm{d}\bm{\theta}^{d}}(\bm{\theta})[\mathbf{u},\cdots,\mathbf{u}]+o(\|\mathbf{u}\|^{D}).$ (4)
The terms of the Taylor expansion contain higher-order information about $\mathcal{L}$ in the direction $\mathbf{u}$. In particular, they can be used to predict how $\mathcal{L}(\bm{\theta})$ would change if $\bm{\theta}$ was translated in the direction of $\mathbf{u}$. Additionally, computing the first $D$ terms has a complexity of order $D\times P$, which is manageable even for large models. The trick that allows for such a low complexity, the Hessian-vector product, was proposed by Pearlmutter (1994) and is recalled in Appendix A.
@@
Now, let us assume that, in the practical implementation of a gradient-based method of optimization of $\mathcal{L}(\bm{\theta})$, $\bm{\theta}$ is represented by a tuple of tensors $(\mathbf{T}^{1},\cdots,\mathbf{T}^{S})$. So, each Taylor term can be expressed as:
$\frac{\mathrm{d}^{d}\mathcal{L}}{\mathrm{d}\bm{\theta}^{d}}(\bm{\theta})[\mathbf{u},\cdot\,\cdot,\mathbf{u}]$ $=\sum_{s_{1}=1}^{S}\cdots\sum_{s_{d}=1}^{S}\frac{\partial^{d}\mathcal{L}}{\partial\mathbf{T}^{s_{1}}\cdots\partial\mathbf{T}^{s_{d}}}(\bm{\theta})[\mathbf{U}^{s_{1}},\cdots,\mathbf{U}^{s_{d}}]=\mathbf{D}_{\bm{\theta}}^{d}(\mathbf{u})[\mathds{1}_{S},\cdots,\mathds{1}_{S}],$ (5)
where $\mathds{1}_{S}\in\mathbb{R}^{S}$ is a vector full of ones, the tuple of tensors $(\mathbf{U}^{1},\cdots,\mathbf{U}^{S})$ represents $\mathbf{u}$,222$(\mathbf{U}^{1},\cdots,\mathbf{U}^{S})$ is to $\mathbf{u}$ as $(\mathbf{T}^{1},\cdots,\mathbf{T}^{S})$ is to $\bm{\theta}$. and $\mathbf{D}_{\bm{\theta}}^{d}(\mathbf{u})\in\mathbb{R}^{S^{d}}$ is a tensor of order $d$ with size $S$ in every dimension with values:
$(\mathbf{D}_{\bm{\theta}}^{d}(\mathbf{u}))_{s_{1},\cdots,s_{d}}$ $=\frac{\partial^{d}\mathcal{L}}{\partial\mathbf{T}^{s_{1}}\cdots\partial\mathbf{T}^{s_{d}}}(\bm{\theta})[\mathbf{U}^{s_{1}},\cdots,\mathbf{U}^{s_{d}}]$ (6)
$=\sum_{i_{1}=1}^{P_{s_{1}}}\cdots\sum_{i_{1}=1}^{P_{s_{d}}}\frac{\partial^{d}\mathcal{L}}{\partial T^{s_{1}}_{i_{1}}\cdots\partial T^{s_{d}}_{i_{d}}}(\bm{\theta})U^{s_{1}}_{i_{1}}\cdots U^{s_{d}}_{i_{d}},$ (7)
where $P_{s}$ is the number of coefficients of the tensor $\mathbf{T}^{s}$. Thus, $\mathbf{D}_{\bm{\theta}}^{d}(\mathbf{u})$ is a tensor of order $d$ and size $S$ in every dimension resulting from a partial contraction of the full derivative $\frac{\mathrm{d}^{d}\mathcal{L}}{\mathrm{d}\bm{\theta}^{d}}(\bm{\theta})$. Moreover, the trick of Pearlmutter (1994) also applies to the computation of $\mathbf{D}_{\bm{\theta}}^{d}(\mathbf{u})$, which is then much less expensive to compute than the full derivative (see Appendix A).
Properties of $\mathbf{D}_{\bm{\theta}}^{d}(\mathbf{u})$.
We show a comparison between the three techniques in Table 1. If $S$ is small enough, computing $\mathbf{D}_{\bm{\theta}}^{d}(\mathbf{u})$ becomes feasible for $d\geq 2$. For usual multilayer perceptrons with $L$ layers, there is one tensor of weights and one vector of biases per layer, so $S=2L$. This allows to compute $\mathbf{D}_{\bm{\theta}}^{d}(\mathbf{u})$ in practice for $d=2$ even when $L\approx 20$.
@@
Full derivative $\frac{\mathrm{d}^{d}\mathcal{L}}{\mathrm{d}\bm{\theta}^{d}}(\bm{\theta})$ $P^{d}$ $P^{d}$
@@
Taylor term $\mathbf{D}_{\bm{\theta}}^{d}(\mathbf{u})[\mathds{1}_{S},\cdots,\mathds{1}_{S}]$ $1$ $d\times P$
Tensor $\mathbf{D}_{\bm{\theta}}^{d}(\mathbf{u})$ $S^{d}$ $S^{d-1}\times P$
@@
According to Eqn. 5, the Taylor term can be obtained by full contraction of $\mathbf{D}_{\bm{\theta}}^{d}(\mathbf{u})$. However, $\mathbf{D}_{\bm{\theta}}^{d}(\mathbf{u})$, is a tensor of size $S^{d}$, and cannot be obtained from the Taylor term, which is only a scalar. Thus, the tensors $\mathbf{D}_{\bm{\theta}}^{d}(\mathbf{u})$ extract more information than the Taylor terms, while keeping a reasonable computational cost. Moreover, their off-diagonal elements give access to information about one-to-one interactions between tensors $(\mathbf{T}^{1},\cdots,\mathbf{T}^{S})$ when they are processed in the function $\mathcal{L}$.
4 A scalable second-order optimization method
4.1 Presentation of the method
@@
The method presented here consists in partitioning the set of indices of parameters $\{1,\cdots,P\}$ into $S$ subsets $(\mathcal{I}_{s})_{1\leq s\leq S}$, assigning for all $1\leq s\leq S$ the same learning rate $\eta_{s}$ to the parameters $(\theta_{p})_{p\in\mathcal{I}_{s}}$, and finding the vector of learning rates $\bm{\eta}=(\eta_{1},\cdots,\eta_{S})$ optimizing the decrease of the loss $\mathcal{L}$ for the current training step $t$, by using its order-2 Taylor approximation. [^3] With the notation of Section 3, $\mathcal{I}_{s}$ is the set of indices $p$ of the parameters $\theta_{p}$ belonging to the tensor $\mathbf{T}^{s}$, so the scalars $(\theta_{p})_{p\in\mathcal{I}_{s}}$ correspond to the scalars belonging to $\mathbf{T}^{s}$. So, everything is as if a specific learning rate $\eta_{s}$ is assigned to each $\mathbf{T}^{s}$. Formally, given a direction $\mathbf{u}_{t}\in\mathbb{R}^{P}$ in the parameter space (typically, $\mathbf{u}_{t}=\mathbf{g}_{t}$, the gradient) and $\mathbf{U}_{t}:=\mathrm{Diag}(\mathbf{u}_{t})\in\mathbb{R}^{P\times P}$, we consider the training step: $\bm{\theta}_{t+1}:=\bm{\theta}_{t}-\mathbf{U}_{t}\mathbf{I}_{P:S}\bm{\eta}_{t}$, that is a training step in a direction based on $\mathbf{u}_{t}$, distorted by a subset-wise step size $\bm{\eta}_{t}$.
Then, we minimize the order-2 Taylor approximation of $\mathcal{L}(\bm{\theta}_{t+1})-\mathcal{L}(\bm{\theta}_{t})$: $\bm{\Delta}_{2}(\bm{\eta}_{t}):=-\mathbf{g}_{t}^{T}\mathbf{U}_{t}\mathbf{I}_{P:S}\bm{\eta}_{t}+\frac{1}{2}\bm{\eta}_{t}^{T}\mathbf{I}_{S:P}\mathbf{U}_{t}\mathbf{H}_{t}\mathbf{U}_{t}\mathbf{I}_{P:S}\bm{\eta}_{t}$, which gives:
$\bm{\theta}_{t+1}$ $=\bm{\theta}_{t}-\mathbf{U}_{t}\mathbf{I}_{P:S}\bm{\eta}_{t}^{*},$ $\bm{\eta}_{t}^{*}$ $:=(\mathbf{I}_{S:P}\mathbf{U}_{t}\mathbf{H}_{t}\mathbf{U}_{t}\mathbf{I}_{P:S})^{-1}\mathbf{I}_{S:P}\mathbf{U}_{t}\mathbf{g}_{t},$ (8)
@@
where $\mathbf{I}_{S:P}\in\mathbb{R}^{S\times P}$ is the partition matrix, verifying $(\mathbf{I}_{S:P})_{sp}=1$ if $p\in\mathcal{I}_{s}$ and 00 otherwise, and $\mathbf{I}_{P:S}:=\mathbf{I}_{S:P}^{T}$. Alternatively, $\bm{\eta}_{t}^{*}$ can be written (details are provided in Appendix B):
@@
With the notation of Section 3, $\bar{\mathbf{H}}_{t}=\mathbf{D}_{\bm{\theta}_{t}}^{(2)}(\mathbf{u}_{t})$ and $\bar{\mathbf{g}}_{t}=\mathbf{D}_{\bm{\theta}_{t}}^{(1)}(\mathbf{u}_{t})$. Incidentally, computing $\bar{\mathbf{H}}$ is of complexity $SP$, and solving the system $\bar{\mathbf{H}}\mathbf{x}=\bar{\mathbf{g}}$ is of complexity $S^{2}$.
4.2 Regularizing $\bar{\mathbf{H}}$ by using order-3 information
The method proposed in Section 4.1 requires to compute $\bm{\eta}^{*}=\bar{\mathbf{H}}^{-1}\bar{\mathbf{g}}$. Usually, inverting such a linear system at every step is considered as hazardous and unstable. Therefore, when using Newton’s method, instead of computing descent direction $\mathbf{u}:=\mathbf{H}^{-1}\mathbf{g}$, it is very common to add a regularization term: $\mathbf{u}_{\lambda}:=\left(\mathbf{H}+\lambda\mathbf{I}\right)^{-1}\mathbf{g}$ (Nocedal & Wright, 1999, Chap. 6.3).
However, the theoretical ground of such a regularization technique is not fully satisfactory. Basically, the main problem is not having a matrix $\bar{\mathbf{H}}$ with close-to-zero eigenvalues: after all, if the loss landscape is very flat in a specific direction, it is better to make a large training step. The problem lies in the order-2 approximation of the loss made in the training step 8, as well as in Newton’s method: instead of optimizing the true decrease of the loss, we optimize the decrease of its order-2 approximation. Thus, the practical question is: does this approximation faithfully model the loss at the current point $\bm{\theta}_{t}$, in a region that also includes the next point $\bm{\theta}_{t+1}$?
To answer this question, one has to take into account order-3 information, and regularize $\bar{\mathbf{H}}$ in such a way that the resulting update remains in a region around $\bm{\theta}_{t}$ where the cubic term of the Taylor approximation is negligible. In practice, we propose an anisotropic version of Nesterov’s cubic regularization (Nesterov & Polyak, 2006).

### v4

@@ 3 Summarizing higher-order information
Let us consider the minimization of a loss function $\mathcal{L}:\mathbb{R}^{P}\rightarrow\mathbb{R}$ according to a variable $\boldsymbol{\theta}\in\mathbb{R}^{P}$.
@@
The order-$d$ derivative of $\mathcal{L}$ at a point $\boldsymbol{\theta}$, that we denote by $\frac{\mathrm{d}^{d}\mathcal{L}}{\mathrm{d}\boldsymbol{\theta}^{d}}(\boldsymbol{\theta})$, can be viewed as either a $d$-linear form (see [7] and Appendix 0.L) or as an order-$d$ tensor belonging to $\mathbb{R}^{P^{d}}$. For convenience, we will use the latter: the coefficients of the tensor $\mathbf{A}=\frac{\mathrm{d}^{d}\mathcal{L}}{\mathrm{d}\boldsymbol{\theta}^{d}}(\boldsymbol{\theta})\in\mathbb{R}^{P^{d}}$ are $A_{i_{1},\cdots,i_{d}}=\frac{\partial^{d}\mathcal{L}}{\partial\theta_{i_{1}}\cdots\partial\theta_{i_{d}}}(\boldsymbol{\theta})$, where $(i_{1},\cdots,i_{d})\in\{1,\cdots,P\}^{d}$ is a multi-index. For a tensor $\mathbf{A}\in\mathbb{R}^{P^{d}}$, we will use the following notation for tensor contraction:
$\forall(\mathbf{u}_{1},\cdots,\mathbf{u}_{d})\in\mathbb{R}^{P}\times\cdots\times\mathbb{R}^{P},$
$\mathbf{A}[\mathbf{u}^{1},\cdots,\mathbf{u}^{d}]:=\sum_{i_{1}=1}^{P}\cdots\sum_{i_{d}=1}^{P}A_{i_{1},\cdots,i_{d}}u^{1}_{i_{1}}\cdots u^{d}_{i_{d}}.$ (3)
The order-$d$ derivative $\frac{\mathrm{d}^{d}\mathcal{L}}{\mathrm{d}\boldsymbol{\theta}^{d}}(\boldsymbol{\theta})\in\mathbb{R}^{P^{d}}$ contains $P^{d}$ scalars. But, even when considering its symmetries, it is computationally too expensive to compute it exactly for $d\geq 2$ in most cases. For instance, it is not even possible to compute numerically the full Hessian of $\mathcal{L}$ according to the parameters of a small neural network, i.e., with $P=10^{5}$ and $d=2$, the Hessian contains $P^{d}=10^{10}$ scalars.
@@
At the opposite, one can obtain cheap higher-order information about $\mathcal{L}$ at $\boldsymbol{\theta}$ by considering a specific direction $\mathbf{u}\in\mathbb{R}^{P}$. The Taylor expansion of $\mathcal{L}(\boldsymbol{\theta}+\mathbf{u})$ gives:
$\mathcal{L}(\boldsymbol{\theta}+\mathbf{u})=\mathcal{L}(\boldsymbol{\theta})+\sum_{d=1}^{D}\frac{1}{d!}\frac{\mathrm{d}^{d}\mathcal{L}}{\mathrm{d}\boldsymbol{\theta}^{d}}(\boldsymbol{\theta})[\mathbf{u},\cdots,\mathbf{u}]+o(\|\mathbf{u}\|^{D}).$ (4)
The terms of the Taylor expansion contain higher-order information about $\mathcal{L}$ in the direction $\mathbf{u}$. In particular, they can be used to predict how $\mathcal{L}(\boldsymbol{\theta})$ would change if $\boldsymbol{\theta}$ was translated in the direction of $\mathbf{u}$. Additionally, computing the first $D$ terms has a complexity of order $D\times P$, which is manageable even for large models. The trick that allows for such a low complexity, the Hessian-vector product, was proposed by [28] and is recalled in Appendix 0.A.
@@
Now, let us assume that, in the practical implementation of a gradient-based method of optimization of $\mathcal{L}(\boldsymbol{\theta})$, $\boldsymbol{\theta}$ is represented by a tuple of tensors $(\mathbf{T}^{1},\cdots,\mathbf{T}^{S})$. So, each Taylor term can be expressed as:
$\frac{\mathrm{d}^{d}\mathcal{L}}{\mathrm{d}\boldsymbol{\theta}^{d}}(\boldsymbol{\theta})[\mathbf{u},\cdot\,\cdot,\mathbf{u}]$ $=\sum_{s_{1}=1}^{S}\cdots\sum_{s_{d}=1}^{S}\frac{\partial^{d}\mathcal{L}}{\partial\mathbf{T}^{s_{1}}\cdots\partial\mathbf{T}^{s_{d}}}(\boldsymbol{\theta})[\mathbf{U}^{s_{1}},\cdots,\mathbf{U}^{s_{d}}]$
$=\mathbf{D}_{\boldsymbol{\theta}}^{d}(\mathbf{u})[\mathds{1}_{S},\cdots,\mathds{1}_{S}],$ (5)
where $\mathds{1}_{S}\in\mathbb{R}^{S}$ is a vector full of ones, the tuple of tensors $(\mathbf{U}^{1},\cdots,\mathbf{U}^{S})$ represents $\mathbf{u}$,222$(\mathbf{U}^{1},\cdots,\mathbf{U}^{S})$ is to $\mathbf{u}$ as $(\mathbf{T}^{1},\cdots,\mathbf{T}^{S})$ is to $\boldsymbol{\theta}$. and $\mathbf{D}_{\boldsymbol{\theta}}^{d}(\mathbf{u})\in\mathbb{R}^{S^{d}}$ is a tensor of order $d$ with size $S$ in every dimension s.t.:
$(\mathbf{D}_{\boldsymbol{\theta}}^{d}(\mathbf{u}))_{s_{1},\cdots,s_{d}}$ $=\frac{\partial^{d}\mathcal{L}}{\partial\mathbf{T}^{s_{1}}\cdots\partial\mathbf{T}^{s_{d}}}(\boldsymbol{\theta})[\mathbf{U}^{s_{1}},\cdots,\mathbf{U}^{s_{d}}]$ (6)
$=\sum_{i_{1}=1}^{P_{s_{1}}}\cdots\sum_{i_{1}=1}^{P_{s_{d}}}\frac{\partial^{d}\mathcal{L}}{\partial T^{s_{1}}_{i_{1}}\cdots\partial T^{s_{d}}_{i_{d}}}(\boldsymbol{\theta})U^{s_{1}}_{i_{1}}\cdots U^{s_{d}}_{i_{d}},$ (7)
where $P_{s}$ is the number of coefficients of the tensor $\mathbf{T}^{s}$. Thus, $\mathbf{D}_{\boldsymbol{\theta}}^{d}(\mathbf{u})$ is a tensor of order $d$ and size $S$ in every dimension resulting from a partial contraction of the full derivative $\frac{\mathrm{d}^{d}\mathcal{L}}{\mathrm{d}\boldsymbol{\theta}^{d}}(\boldsymbol{\theta})$. Moreover, the trick of [28] also applies to the computation of $\mathbf{D}_{\boldsymbol{\theta}}^{d}(\mathbf{u})$, which is then much less expensive to compute than the full derivative (see Appendix 0.A).
Properties of $\mathbf{D}_{\boldsymbol{\theta}}^{d}(\mathbf{u})$.
We show a comparison between the three techniques in Table 1. If $S$ is small enough, computing $\mathbf{D}_{\boldsymbol{\theta}}^{d}(\mathbf{u})$ becomes feasible for $d\geq 2$. For usual multilayer perceptrons with $L$ layers, there is one tensor of weights and one vector of biases per layer, so $S=2L$. This allows to compute $\mathbf{D}_{\boldsymbol{\theta}}^{d}(\mathbf{u})$ in practice for $d=2$ even when $L\approx 20$.
@@
Full derivative $\frac{\mathrm{d}^{d}\mathcal{L}}{\mathrm{d}\boldsymbol{\theta}^{d}}(\boldsymbol{\theta})$ $P^{d}$ $P^{d}$
@@
Taylor term $\mathbf{D}_{\boldsymbol{\theta}}^{d}(\mathbf{u})[\mathds{1}_{S},\cdots,\mathds{1}_{S}]$ $1$ $d\times P$
Tensor $\mathbf{D}_{\boldsymbol{\theta}}^{d}(\mathbf{u})$ $S^{d}$ $S^{d-1}\times P$
@@
According to Eqn. (5), the Taylor term can be obtained by full contraction of $\mathbf{D}_{\boldsymbol{\theta}}^{d}(\mathbf{u})$. However, $\mathbf{D}_{\boldsymbol{\theta}}^{d}(\mathbf{u})$, is a tensor of size $S^{d}$, and cannot be obtained from the Taylor term, which is only a scalar. Thus, the tensors $\mathbf{D}_{\boldsymbol{\theta}}^{d}(\mathbf{u})$ extract more information than the Taylor terms, while keeping a reasonable computational cost. Moreover, their off-diagonal elements give access to information about one-to-one interactions between tensors $(\mathbf{T}^{1},\cdots,\mathbf{T}^{S})$ when they are processed in the function $\mathcal{L}$.
4 Application: computing per-layer learning rates
@@
To build per-layer (or per-subset-of-parameters) learning rates, we partition the set of indices of parameters $\{1,\cdots,P\}$ into $S$ subsets $(\mathcal{I}_{s})_{1\leq s\leq S}$, we assign for all $1\leq s\leq S$ the same learning rate $\eta_{s}$ to the parameters $(\theta_{p})_{p\in\mathcal{I}_{s}}$, and we find the vector of learning rates $\boldsymbol{\eta}=(\eta_{1},\cdots,\eta_{S})$ optimizing the decrease of the loss $\mathcal{L}$ for the current training step $t$, by using its order-2 Taylor approximation. [^3] With the notation of Section 3, $\mathcal{I}_{s}$ is the set of indices $p$ of the parameters $\theta_{p}$ belonging to the tensor $\mathbf{T}^{s}$, so the scalars $(\theta_{p})_{p\in\mathcal{I}_{s}}$ correspond to the scalars belonging to $\mathbf{T}^{s}$. So, everything is as if a specific learning rate $\eta_{s}$ is assigned to each $\mathbf{T}^{s}$. Formally, given a direction $\mathbf{u}_{t}\in\mathbb{R}^{P}$ in the parameter space (typically, $\mathbf{u}_{t}=\mathbf{g}_{t}$, the gradient) and $\mathbf{U}_{t}:=\mathrm{Diag}(\mathbf{u}_{t})\in\mathbb{R}^{P\times P}$, we consider the training step: $\boldsymbol{\theta}_{t+1}:=\boldsymbol{\theta}_{t}-\mathbf{U}_{t}\mathbf{I}_{P:S}\boldsymbol{\eta}_{t}$, that is a training step in a direction based on $\mathbf{u}_{t}$, distorted by a subset-wise step size $\boldsymbol{\eta}_{t}$. Then, we minimize the order-2 Taylor approximation of $\mathcal{L}(\boldsymbol{\theta}_{t+1})-\mathcal{L}(\boldsymbol{\theta}_{t})$: $\boldsymbol{\Delta}_{2}(\boldsymbol{\eta}_{t}):=-\mathbf{g}_{t}^{T}\mathbf{U}_{t}\mathbf{I}_{P:S}\boldsymbol{\eta}_{t}+\frac{1}{2}\boldsymbol{\eta}_{t}^{T}\mathbf{I}_{S:P}\mathbf{U}_{t}\mathbf{H}_{t}\mathbf{U}_{t}\mathbf{I}_{P:S}\boldsymbol{\eta}_{t}$, which gives:
$\boldsymbol{\theta}_{t+1}$ $=\boldsymbol{\theta}_{t}-\mathbf{U}_{t}\mathbf{I}_{P:S}\boldsymbol{\eta}_{t}^{*},$ $\boldsymbol{\eta}_{t}^{*}$ $:=(\mathbf{I}_{S:P}\mathbf{U}_{t}\mathbf{H}_{t}\mathbf{U}_{t}\mathbf{I}_{P:S})^{-1}\mathbf{I}_{S:P}\mathbf{U}_{t}\mathbf{g}_{t},$ (8)
@@
where $\mathbf{I}_{S:P}\in\mathbb{R}^{S\times P}$ is the partition matrix, verifying $(\mathbf{I}_{S:P})_{sp}=1$ if $p\in\mathcal{I}_{s}$ and 00 otherwise, and $\mathbf{I}_{P:S}:=\mathbf{I}_{S:P}^{T}$. Alternatively, $\boldsymbol{\eta}_{t}^{*}$ can be written (details are provided in Appendix 0.B):
@@
With the notation of Section 3, $\bar{\mathbf{H}}_{t}=\mathbf{D}_{\boldsymbol{\theta}_{t}}^{(2)}(\mathbf{u}_{t})$ and $\bar{\mathbf{g}}_{t}=\mathbf{D}_{\boldsymbol{\theta}_{t}}^{(1)}(\mathbf{u}_{t})$. Incidentally, computing $\bar{\mathbf{H}}$ is of complexity $SP$, and solving $\bar{\mathbf{H}}\mathbf{x}=\bar{\mathbf{g}}$ is of complexity $S^{2}$.
5 Application: optimization method
5.1 Presentation
Now that we can can compute per-layer learning rates, we decide to incorporate them into an optimization method. However, computing them requires to compute $\bar{\mathbf{H}}^{-1}\bar{\mathbf{g}}$. Usually, inverting such a linear system at every step is considered as hazardous and unstable. Therefore, when using Newton’s method, instead of computing descent direction $\mathbf{u}:=\mathbf{H}^{-1}\mathbf{g}$, it is very common to add a regularization term: $\mathbf{u}_{\lambda}:=\left(\mathbf{H}+\lambda\mathbf{I}\right)^{-1}\mathbf{g}$ [25, Chap. 6.3].
However, the theoretical ground of such a regularization technique is not fully satisfactory. Basically, the main problem is not having a matrix $\bar{\mathbf{H}}$ with close-to-zero eigenvalues: after all, if the loss landscape is very flat in a specific direction, it is better to make a large training step. The problem lies in the order-2 approximation of the loss made in the training step (8), as well as in Newton’s method: instead of optimizing the true decrease of the loss, we optimize the decrease of its order-2 approximation. Thus, the practical question is: does this approximation faithfully model the loss at the current point $\boldsymbol{\theta}_{t}$, in a region that also includes the next point $\boldsymbol{\theta}_{t+1}$?
To answer this question, one has to take into account order-3 information, and regularize $\bar{\mathbf{H}}$ so that the resulting update remains in a region around $\boldsymbol{\theta}_{t}$ where the cubic term of the Taylor approximation is negligible. In practice, we propose an anisotropic version of Nesterov’s cubic regularization [24].